Find the general solution of the following differential equation
3y" +2y' +y = 0
1
Expert's answer
2021-09-22T17:26:17-0400

Corresponding (auxiliary) equation


"3r^2+2r+1=0"

"D=(2)^2-4(3)(1)=-8"

"r=\dfrac{-2\pm\sqrt{-8}}{2(3)}=\dfrac{-1\pm\sqrt{2}}{3}"

"r_1=\dfrac{-1-\sqrt{2}}{3},r_2=\dfrac{-1+\sqrt{2}}{3}"

The general solution of the given differential equation is


"y(t)=c_1e^{-{1 \over 3}t}\cos({\sqrt{2} \over 3}t)+c_2e^{-{1 \over 3}t}\sin({\sqrt{2} \over 3}t)"